OpenStudy (anonymous):

a juggler tosses 3 talls alternately vertically upward . Each ball has an initial velocity of 5m/s . How high does each ball rise ? How long does each ball remain in the air?

OpenStudy (anonymous):

v = vo + a t a = -9.8 m/s^2 0 = 5m/s - [9.8 m/s^2] t t = 5/9.8 = 0.51 s on the way up and another 0.51 sec on the way back down height: m g h = P.E. = original K.E = 0.5 m vo^2 h = 0.5 vo^2 / g = 0.5 (25)/(9.8) = 1.27 check: s = vo t - 0.5 g t^2 = 5(.51) - 0.5(9.8) (.51)^2 s= 1.27 yes Actually, number may have been chosen to let you use g=10m/s^2. etc.